Calculate the acceleration due to gravity at a depth 384 km from the surface of the earth


1
Expert's answer
2021-08-07T12:37:38-0400

The acceleration due to gravity is given by

"g=G\frac{M}{R^2}=G\frac{M}{(R_{\rm Earth}-h)^2}"

"M=\rho V=\frac{4}{3}\pi\rho (R_{\rm Earth}-h)^3"

"g=\frac{4}{3}\pi\rho G (R_{\rm Earth}-h)\\"

"=\frac{4}{3}*3.14*5510*6.67*10^{-11}*(6378-384)*10^3\\=9.2\:\rm m/s^2"